OpenStudy (anonymous):

Equation R: y = 5x + 10 Equation S: y = 5x + 5 Which of the following options is true about the solution to the given set of equations? A. one B. two C. infinite D. none

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

hi do u need help

OpenStudy (anonymous):

yess @Hike_The_Gamer

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

but not 100% sure?

OpenStudy (anonymous):

ok can you help me with another @Hike_The_Gamer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

neee

OpenStudy (anonymous):

Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? A. x=2x+1 B. x+9=2x C. x+1=2x+9 D. x+9=2x+1

OpenStudy (anonymous):

@Hike_The_Gamer

OpenStudy (anonymous):

on a postive note it not c and a

OpenStudy (anonymous):

what do u think

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

b

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

thank you and yes hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y = –2x + 1 6x + 2y = 22 Which of the following steps could be used to solve by substitution? A. 6x+2(-2x+1)=22 B. -2x+1=6x+2y C. 6(-2x+1+2y=22 D. 6(y=-2x+1)

OpenStudy (anonymous):

its esay

OpenStudy (anonymous):

really

OpenStudy (anonymous):

yeah its a do u have more i can do 1 more only

OpenStudy (anonymous):

thanks and ok

OpenStudy (anonymous):

u have one more

OpenStudy (anonymous):

no thats is thanks

OpenStudy (anonymous):

bye

Directrix (directrix):

@Hike_The_Gamer Do not post direct answers with no solution of how to get that answer. Your direct answers were not all corrrect. http://openstudy.com/code-of-conduct